Suppose that your brother wants to have $119000 to buy video equipment in five years. If he can invest his money now at 12.5% compounded continuously, then how much must he invest today? Round your answer to the nearest dollar. Do not use a dollar sign or a comma
Answer by Theo(13342)

the formula for continuous compounding is f = p * e^(rt)

f is the future value.
p is the present value.
r is the interest rate per time period.
t is the number of time periods.

in your problem, the equation becomes 119000 = p * e^(.125*5)

solve for p to get p = 119000 / (e^(.125*5))

result is p = 63696.10999

round to nearest dollar to get p = 63696